# second generation 3G third generation

### v1

@@ III-B TUCKER-Based Receiver
    Figure 4: TUCKER tensor model: In this scenario, the tensor from $i$-th frame, $\boldsymbol{\mathcal{Y}}_{i}$ in $(18)$, is rearranged to derive a fourth-order TUCKER model for the received signal tensor.
@@
    $\boldsymbol{Y}_{i}=\Big(\boldsymbol{X}\otimes\boldsymbol{HS}\Big)\text{diag}(\boldsymbol{g}_{i})\boldsymbol{\Psi}^{\text{T}}\;,$ (34)
@@
    (1) $=\Big[\boldsymbol{Y}_{1}\;\boldsymbol{Y}_{2}\;\dots\;\boldsymbol{Y}_{I}\Big]$ (35)
    $=\Big(\boldsymbol{X}\otimes\boldsymbol{HS}\Big)\Big(\overline{\boldsymbol{G}}\diamond\boldsymbol{\Psi}\Big)^{\text{T}}\;\in\mathbb{C}^{TM_{R}\times IK}\;.$
    Note that the equation (35) can be see as a third order tensor $\widetilde{\boldsymbol{\mathcal{Y}}}\in\mathbb{C}^{M_{R}\times T\times IK}$ that corresponds to a TUCKER model where the its $n$-mode product is
    $\widetilde{\boldsymbol{\mathcal{Y}}}=(\overline{\boldsymbol{\mathcal{G}}}\diamond_{3}\boldsymbol{\mathcal{T}})\times_{1}\boldsymbol{HS}\times_{2}\boldsymbol{X}\times_{3}\boldsymbol{I}_{IK}\;\in\mathbb{C}^{M_{R}\times T\times IK}\;,$ (36)
    where the core tensor $\boldsymbol{\mathcal{C}}=\Big(\overline{\boldsymbol{\mathcal{G}}}\diamond_{3}\boldsymbol{\mathcal{T}}\Big)\in\mathbb{C}^{N\times M_{T}\times IK}$ has a special structure which is given by $3$-mode Khatri-Rao product between two tensors, which are
    $\overline{\boldsymbol{\mathcal{G}}}\doteq\Big[\overline{\boldsymbol{G}}_{1}\sqcup_{3}\;\overline{\boldsymbol{G}}_{2}\sqcup_{3}\;\dots\;\sqcup_{3}\overline{\boldsymbol{G}_{I}}\Big]\;\mathbb{C}^{N\times M_{T}\times I}$ (37)
@@
    Note that the channel $\overline{\boldsymbol{G}}$ lies within the core tensor. This implies that, in a straightforward approach, one would need at least a two-stage procedure: first estimate the core tensor, then extract $\overline{\boldsymbol{G}}$. Instead of reducing the tensor order, however, we can preserve the fourth-order structure and reinterpret it as a TUCKER model. This can be done by applying the multimode unfolding concept, which reveals that the PARATUCK-(2,4) model can be expressed as a fourth-order TUCKER tensor model. This relationship is depicted in Fig. 4, where the $I$ tensors $\boldsymbol{\mathcal{Y}}_{i}$ ($i=1,\ldots,I$) are transformed, via multimode unfolding, into a TUCKER tensor $\boldsymbol{\mathcal{Q}}$ (in tensor network notation). This transformation preserves the original dimensions but changes the underlying tensor model. Such a reformulation enables the derivation of a single-stage solution, as discussed next.
@@
    We know that given $\boldsymbol{A}\in\mathbb{C}^{N_{A}\times L}$ and $\boldsymbol{B}\in\mathbb{C}^{N_{B}\times L}$, the Khatri-Rao product between them can be expressed as $\boldsymbol{A}\diamond\boldsymbol{B}=(\boldsymbol{A}\otimes\boldsymbol{B})\boldsymbol{\Xi}\in\mathbb{C}^{N_{A}N_{B}\times L}$ where $\boldsymbol{\Xi}$ is an $L^{2}\times L$ selection matrix that extracts the column-wise Kronecker product. Now, applying this Khatri-Rao and Kronecker product relationship to the $1$-mode unfolding in (35), we can rewrite it as
    (1) $=\Big(\boldsymbol{X}\otimes\boldsymbol{HS}\Big)\Big(\overline{\boldsymbol{G}}\diamond\boldsymbol{\Psi}\Big)^{\text{T}}\;\in\mathbb{C}^{TM_{R}\times IK}\;$ (39)
    $=\Big(\boldsymbol{X}\otimes\boldsymbol{HS}\Big)\Big((\overline{\boldsymbol{G}}\otimes\boldsymbol{\Psi})\boldsymbol{\Xi}\Big)^{\text{T}}$
    $=\Big(\boldsymbol{X}\otimes\boldsymbol{HS}\Big)\boldsymbol{\Xi}^{\text{T}}\Big(\overline{\boldsymbol{G}}\otimes\boldsymbol{\Psi}\Big)^{\text{T}},$
@@
    where $\boldsymbol{\Xi}\in\mathbb{C}^{(M_{T}N)^{2}\times M_{T}N}$ is the selection matrix such that $\boldsymbol{\Xi}=\boldsymbol{I}_{M_{T}N}\diamond\boldsymbol{I}_{M_{T}N}$. From these observation, we can note that $\boldsymbol{\mathcal{Y}}]_{(1)}$ can be see as generalized unfolding of a tensor $\boldsymbol{\mathcal{Q}}$, this is $\boldsymbol{\mathcal{Q}}]_{([1,2],[3,4])}=[\boldsymbol{\mathcal{Y}}]_{(1)}$. Then,
    ([1,2],[3,4]) $=\Big(\boldsymbol{X}\otimes\boldsymbol{HS}\Big)(\boldsymbol{I}_{M_{T}N}\diamond\boldsymbol{I}_{M_{T}N})^{\text{T}}\Big(\overline{\boldsymbol{G}}\otimes\boldsymbol{\Psi}\Big)^{\text{T}},$ (40)
@@
    where $(\boldsymbol{I}_{M_{T}N}\diamond\boldsymbol{I}_{M_{T}N})^{\text{T}}=[\boldsymbol{\mathcal{C}}]_{([1,2],[3,4])}\in\mathbb{C}^{M_{T}N\times(M_{T}N)^{2}}$ is the multimode unfolding of a fourth order core tensor $\boldsymbol{\mathcal{C}}\in\mathbb{C}^{N\times M_{T}\times M_{T}N\times M_{T}N}$ while the equation (40) denotes the multimode unfolding of the tensor $\boldsymbol{\mathcal{Q}}\in\mathbb{C}^{M_{R}\times T\times K\times I}$, whose the $n$-mode product notation is
    $\boldsymbol{\mathcal{Q}}=\boldsymbol{\mathcal{C}}\times_{1}\boldsymbol{HS}\times_{2}\boldsymbol{X}\times_{3}\boldsymbol{\Psi}\times_{4}\overline{\boldsymbol{G}}.$ (41)
@@
    The fourth-order tensor $\boldsymbol{\mathcal{Q}}$ fits a TUCKER model, where the matrix $\boldsymbol{\Psi}$ is known. To estimate the matrices $\boldsymbol{HS}$, $\boldsymbol{X}$ and $\overline{\boldsymbol{G}}$ it is necessary to solve the following optimization problem:
    $\underset{(\boldsymbol{HS},\boldsymbol{X},\overline{\boldsymbol{G}})}{\arg\min}\|\boldsymbol{\mathcal{Q}}-\boldsymbol{\mathcal{C}}\times_{1}\boldsymbol{HS}\times_{2}\boldsymbol{X}\times_{3}\boldsymbol{\Psi}\times_{4}\overline{\boldsymbol{G}}\|_{F}^{2}.$ (42)
@@
    $\boldsymbol{\mathcal{Q}}]_{(1)}=\boldsymbol{HS}[\boldsymbol{\mathcal{C}}]_{(1)}\big(\overline{\boldsymbol{G}}\otimes\boldsymbol{\Psi}\otimes\boldsymbol{X}\big)^{\text{T}}\in\mathbb{C}^{M_{R}\times IKT},$ (43)
    $\boldsymbol{\mathcal{Q}}]_{(2)}=\boldsymbol{X}[\boldsymbol{\mathcal{C}}]_{(2)}\big(\overline{\boldsymbol{G}}\otimes\boldsymbol{\Psi}\otimes\boldsymbol{HS}\big)^{\text{T}}\in\mathbb{C}^{T\times IKM_{R}},$ (44)
    $\boldsymbol{\mathcal{Q}}]_{(3)}=\boldsymbol{\Psi}[\boldsymbol{\mathcal{C}}]_{(3)}\big(\overline{\boldsymbol{G}}\otimes\boldsymbol{X}\otimes\boldsymbol{HS}\big)^{\text{T}}\in\mathbb{C}^{K\times ITM_{R}},$ (45)
    $\boldsymbol{\mathcal{Q}}]_{(4)}=\overline{\boldsymbol{G}}[\boldsymbol{\mathcal{C}}]_{(4)}\big(\boldsymbol{\Psi}\otimes\boldsymbol{X}\otimes\boldsymbol{HS}\big)^{\text{T}}\in\mathbb{C}^{I\times KTM_{R}},$ (46)
    with $\boldsymbol{\mathcal{C}}]_{(1)}\in\mathbb{C}^{N\times(M_{T}N)^{2}M_{T}}$, $\boldsymbol{\mathcal{C}}]_{(2)}\in\mathbb{C}^{M_{T}\times(M_{T}N)^{2}N}$, $\boldsymbol{\mathcal{C}}]_{(3)}\in\mathbb{C}^{M_{T}N\times(M_{T}N)^{2}}$ and $\boldsymbol{\mathcal{C}}]_{(4)}\in\mathbb{C}^{M_{T}N\times(M_{T}N)^{2}}$. From the first, second, and fourth mode unfoldings, we can derive three sub-optimization linear problems as given
    $\underset{(\boldsymbol{HS})}{\arg\min}\|[\boldsymbol{\mathcal{Q}}]_{(1)}-\boldsymbol{HS}[\boldsymbol{\mathcal{C}}]_{(1)}\big(\overline{\boldsymbol{G}}\otimes\boldsymbol{\Psi}\otimes\boldsymbol{X}\big)^{\text{T}}\|_{F}^{2},$ (47)
    $\underset{(\boldsymbol{X})}{\arg\min}\|[\boldsymbol{\mathcal{Q}}]_{(2)}-\boldsymbol{X}[\boldsymbol{\mathcal{C}}]_{(2)}\big(\overline{\boldsymbol{G}}\otimes\boldsymbol{\Psi}\otimes\boldsymbol{HS}\big)^{\text{T}}\|_{F}^{2},$ (48)
    $\underset{(\overline{\boldsymbol{G}})}{\arg\min}\|[\boldsymbol{\mathcal{Q}}]_{(4)}-\overline{\boldsymbol{G}}[\boldsymbol{\mathcal{C}}]_{(4)}\big(\boldsymbol{\Psi}\otimes\boldsymbol{X}\otimes\boldsymbol{HS}\big)^{\text{T}}\|_{F}^{2}.$ (49)
    To estimate $\boldsymbol{HS}$, $\boldsymbol{X}$ and $\overline{\boldsymbol{G}}$, we solve alternately the LS problems in (47), (48) and (49), respectively, whose the solutions are given by:
    $\hat{\boldsymbol{HS}}=[\boldsymbol{\mathcal{Q}}]_{(1)}\left[[\boldsymbol{\mathcal{C}}]_{(1)}\big(\overline{\boldsymbol{G}}\otimes\boldsymbol{\Psi}\otimes\boldsymbol{X}\big)^{\text{T}}\right]^{\dagger}$ (50)
    $\hat{\boldsymbol{X}}=[\boldsymbol{\mathcal{Q}}]_{(2)}\left[[\boldsymbol{\mathcal{C}}]_{(2)}\big(\overline{\boldsymbol{G}}\otimes\boldsymbol{\Psi}\otimes\boldsymbol{HS}\big)^{\text{T}}\right]^{\dagger}$ (51)
    $\hat{\overline{\boldsymbol{G}}}=[\boldsymbol{\mathcal{Q}}]_{(4)}\left[[\boldsymbol{\mathcal{C}}]_{(4)}\big(\boldsymbol{\Psi}\otimes\boldsymbol{X}\otimes\boldsymbol{HS}\big)^{\text{T}}\right]^{\dagger}.$ (52)
@@
    input : Received signal $\boldsymbol{\mathcal{Q}}$
    output : $\hat{\boldsymbol{H}}$, $\hat{\overline{\boldsymbol{G}}}$ and $\hat{\boldsymbol{X}}$
    1exto $i=0$ initialize $\hat{\overline{\boldsymbol{G}}}_{(i=0)}$ and $\hat{\boldsymbol{X}}_{(i=0)}$ begin
@@
      Find a least squares estimate of $\hat{\boldsymbol{HS}}_{i}$:
    • $\hat{(\boldsymbol{HS})}_{i}=[\boldsymbol{\mathcal{Q}}]_{(1)}\big[\underbrace{[\boldsymbol{\mathcal{C}}]_{(1)}\big(\overline{\boldsymbol{G}}_{(i-1)}\otimes\boldsymbol{\Psi}\otimes\boldsymbol{X}_{(i-1)}\big)^{\text{T}}}_{N\times IKT}\big]^{\dagger}$
@@
    Find a least squares estimate of $\hat{\boldsymbol{X}}_{i}$:
    • $\hat{\boldsymbol{X}}_{i}=[\boldsymbol{\mathcal{Q}}]_{(2)}\big[\underbrace{[\boldsymbol{\mathcal{C}}]_{(2)}\big(\overline{\boldsymbol{G}}_{(i-1)}\otimes\boldsymbol{\Psi}\otimes(\boldsymbol{HS})_{i}\big)^{\text{T}}}_{M_{T}\times IKM_{R}}\big]^{\dagger}$
@@
    Find a least squares estimate of $\hat{\overline{\boldsymbol{G}}}_{i}$:
    • $\hat{\overline{\boldsymbol{G}}}_{i}=[\boldsymbol{\mathcal{Q}}]_{(4)}\big[\underbrace{[\boldsymbol{\mathcal{C}}]_{(4)}\big(\boldsymbol{\Psi}\otimes\boldsymbol{X}_{i}\otimes(\boldsymbol{HS})_{i}\big)^{\text{T}}}_{M_{T}N\times KTM_{R}}\big]^{\dagger}$
@@
    Repeat steps $1$ to $3$ until convergence.
    end
@@
    From (45), an alternative (closed-form) approach can be used to jointly estimate the involved channels $\boldsymbol{H}$, $\boldsymbol{G}$ as well as the data symbol matrix $\boldsymbol{X}$. This approach left-filters the $3$-mode unfolding of the received signal using $(\boldsymbol{\Psi}[\boldsymbol{\mathcal{C}}]_{(1)})^{\dagger}$ followed by solving a “triple-Kronecker” factorization problem o the filtered signal. The algorithm is similar to the one proposed in [SOKAL_Elsevier]. Despite being a less complex estimation method, this approach imposes a more restrictive restriction on the data block length, requiring $K\geq(M_{T}N)^{2}$.

### v2

@@ III-B TUCKER-Based Receiver
    Figure 4: TUCKER tensor model: In this scenario, the tensor from $i$-th frame, $\bm{\mathcal{Y}}_{i}$ in $(18)$, is rearranged to derive a fourth-order TUCKER model for the received signal tensor.
@@
    $\bm{Y}_{i}=\Big(\bm{X}\otimes\bm{HS}\Big)\text{diag}(\bm{g}_{i})\bm{\Psi}^{\text{T}}\;,$ (34)
@@
    (1) $=\Big[\bm{Y}_{1}\;\bm{Y}_{2}\;\dots\;\bm{Y}_{I}\Big]$ (35)
    $=\Big(\bm{X}\otimes\bm{HS}\Big)\Big(\overline{\bm{G}}\diamond\bm{\Psi}\Big)^{\text{T}}\;\in\mathbb{C}^{TM_{R}\times IK}\;.$
    Note that the equation (35) can be see as a third order tensor $\widetilde{\bm{\mathcal{Y}}}\in\mathbb{C}^{M_{R}\times T\times IK}$ that corresponds to a TUCKER model where the its $n$-mode product is
    $\widetilde{\bm{\mathcal{Y}}}=(\overline{\bm{\mathcal{G}}}\diamond_{3}\bm{\mathcal{T}})\times_{1}\bm{HS}\times_{2}\bm{X}\times_{3}\bm{I}_{IK}\;\in\mathbb{C}^{M_{R}\times T\times IK}\;,$ (36)
    where the core tensor $\bm{\mathcal{C}}=\Big(\overline{\bm{\mathcal{G}}}\diamond_{3}\bm{\mathcal{T}}\Big)\in\mathbb{C}^{N\times M_{T}\times IK}$ has a special structure which is given by $3$-mode Khatri-Rao product between two tensors, which are
    $\overline{\bm{\mathcal{G}}}\doteq\Big[\overline{\bm{G}}_{1}\sqcup_{3}\;\overline{\bm{G}}_{2}\sqcup_{3}\;\dots\;\sqcup_{3}\overline{\bm{G}_{I}}\Big]\;\mathbb{C}^{N\times M_{T}\times I}$ (37)
@@
    Note that the channel $\overline{\bm{G}}$ lies within the core tensor. This implies that, in a straightforward approach, one would need at least a two-stage procedure: first estimate the core tensor, then extract $\overline{\bm{G}}$. Instead of reducing the tensor order, however, we can preserve the fourth-order structure and reinterpret it as a TUCKER model. This can be done by applying the multimode unfolding concept, which reveals that the PARATUCK-(2,4) model can be expressed as a fourth-order TUCKER tensor model. This relationship is depicted in Fig. 4, where the $I$ tensors $\bm{\mathcal{Y}}_{i}$ ($i=1,\ldots,I$) are transformed, via multimode unfolding, into a TUCKER tensor $\bm{\mathcal{Q}}$ (in tensor network notation). This transformation preserves the original dimensions but changes the underlying tensor model. Such a reformulation enables the derivation of a single-stage solution, as discussed next.
@@
    We know that given $\bm{A}\in\mathbb{C}^{N_{A}\times L}$ and $\bm{B}\in\mathbb{C}^{N_{B}\times L}$, the Khatri-Rao product between them can be expressed as $\bm{A}\diamond\bm{B}=(\bm{A}\otimes\bm{B})\bm{\Xi}\in\mathbb{C}^{N_{A}N_{B}\times L}$ where $\bm{\Xi}$ is an $L^{2}\times L$ selection matrix that extracts the column-wise Kronecker product. Now, applying this Khatri-Rao and Kronecker product relationship to the $1$-mode unfolding in (35), we can rewrite it as
    (1) $=\Big(\bm{X}\otimes\bm{HS}\Big)\Big(\overline{\bm{G}}\diamond\bm{\Psi}\Big)^{\text{T}}\;\in\mathbb{C}^{TM_{R}\times IK}\;$ (39)
    $=\Big(\bm{X}\otimes\bm{HS}\Big)\Big((\overline{\bm{G}}\otimes\bm{\Psi})\bm{\Xi}\Big)^{\text{T}}$
    $=\Big(\bm{X}\otimes\bm{HS}\Big)\bm{\Xi}^{\text{T}}\Big(\overline{\bm{G}}\otimes\bm{\Psi}\Big)^{\text{T}},$
@@
    where $\bm{\Xi}\in\mathbb{C}^{(M_{T}N)^{2}\times M_{T}N}$ is the selection matrix such that $\bm{\Xi}=\bm{I}_{M_{T}N}\diamond\bm{I}_{M_{T}N}$. From these observation, we can note that $\bm{\mathcal{Y}}]_{(1)}$ can be see as generalized unfolding of a tensor $\bm{\mathcal{Q}}$, this is $\bm{\mathcal{Q}}]_{([1,2],[3,4])}=[\bm{\mathcal{Y}}]_{(1)}$. Then,
    ([1,2],[3,4]) $=\Big(\bm{X}\otimes\bm{HS}\Big)(\bm{I}_{M_{T}N}\diamond\bm{I}_{M_{T}N})^{\text{T}}\Big(\overline{\bm{G}}\otimes\bm{\Psi}\Big)^{\text{T}},$ (40)
@@
    where $(\bm{I}_{M_{T}N}\diamond\bm{I}_{M_{T}N})^{\text{T}}=[\bm{\mathcal{C}}]_{([1,2],[3,4])}\in\mathbb{C}^{M_{T}N\times(M_{T}N)^{2}}$ is the multimode unfolding of a fourth order core tensor $\bm{\mathcal{C}}\in\mathbb{C}^{N\times M_{T}\times M_{T}N\times M_{T}N}$ while the equation (40) denotes the multimode unfolding of the tensor $\bm{\mathcal{Q}}\in\mathbb{C}^{M_{R}\times T\times K\times I}$, whose the $n$-mode product notation is
    $\bm{\mathcal{Q}}=\bm{\mathcal{C}}\times_{1}\bm{HS}\times_{2}\bm{X}\times_{3}\bm{\Psi}\times_{4}\overline{\bm{G}}.$ (41)
@@
    The fourth-order tensor $\bm{\mathcal{Q}}$ fits a TUCKER model, where the matrix $\bm{\Psi}$ is known. To estimate the matrices $\bm{HS}$, $\bm{X}$ and $\overline{\bm{G}}$ it is necessary to solve the following optimization problem:
    $\underset{(\bm{HS},\bm{X},\overline{\bm{G}})}{\arg\min}\|\bm{\mathcal{Q}}-\bm{\mathcal{C}}\times_{1}\bm{HS}\times_{2}\bm{X}\times_{3}\bm{\Psi}\times_{4}\overline{\bm{G}}\|_{F}^{2}.$ (42)
@@
    $\bm{\mathcal{Q}}]_{(1)}=\bm{HS}[\bm{\mathcal{C}}]_{(1)}\big(\overline{\bm{G}}\otimes\bm{\Psi}\otimes\bm{X}\big)^{\text{T}}\in\mathbb{C}^{M_{R}\times IKT},$ (43)
    $\bm{\mathcal{Q}}]_{(2)}=\bm{X}[\bm{\mathcal{C}}]_{(2)}\big(\overline{\bm{G}}\otimes\bm{\Psi}\otimes\bm{HS}\big)^{\text{T}}\in\mathbb{C}^{T\times IKM_{R}},$ (44)
    $\bm{\mathcal{Q}}]_{(3)}=\bm{\Psi}[\bm{\mathcal{C}}]_{(3)}\big(\overline{\bm{G}}\otimes\bm{X}\otimes\bm{HS}\big)^{\text{T}}\in\mathbb{C}^{K\times ITM_{R}},$ (45)
    $\bm{\mathcal{Q}}]_{(4)}=\overline{\bm{G}}[\bm{\mathcal{C}}]_{(4)}\big(\bm{\Psi}\otimes\bm{X}\otimes\bm{HS}\big)^{\text{T}}\in\mathbb{C}^{I\times KTM_{R}},$ (46)
    with $\bm{\mathcal{C}}]_{(1)}\in\mathbb{C}^{N\times(M_{T}N)^{2}M_{T}}$, $\bm{\mathcal{C}}]_{(2)}\in\mathbb{C}^{M_{T}\times(M_{T}N)^{2}N}$, $\bm{\mathcal{C}}]_{(3)}\in\mathbb{C}^{M_{T}N\times(M_{T}N)^{2}}$ and $\bm{\mathcal{C}}]_{(4)}\in\mathbb{C}^{M_{T}N\times(M_{T}N)^{2}}$. From the first, second, and fourth mode unfoldings, we can derive three sub-optimization linear problems as given
    $\underset{(\bm{HS})}{\arg\min}\|[\bm{\mathcal{Q}}]_{(1)}-\bm{HS}[\bm{\mathcal{C}}]_{(1)}\big(\overline{\bm{G}}\otimes\bm{\Psi}\otimes\bm{X}\big)^{\text{T}}\|_{F}^{2},$ (47)
    $\underset{(\bm{X})}{\arg\min}\|[\bm{\mathcal{Q}}]_{(2)}-\bm{X}[\bm{\mathcal{C}}]_{(2)}\big(\overline{\bm{G}}\otimes\bm{\Psi}\otimes\bm{HS}\big)^{\text{T}}\|_{F}^{2},$ (48)
    $\underset{(\overline{\bm{G}})}{\arg\min}\|[\bm{\mathcal{Q}}]_{(4)}-\overline{\bm{G}}[\bm{\mathcal{C}}]_{(4)}\big(\bm{\Psi}\otimes\bm{X}\otimes\bm{HS}\big)^{\text{T}}\|_{F}^{2}.$ (49)
    To estimate $\bm{HS}$, $\bm{X}$ and $\overline{\bm{G}}$, we solve alternately the LS problems in (47), (48) and (49), respectively, whose the solutions are given by:
    $\hat{\bm{HS}}=[\bm{\mathcal{Q}}]_{(1)}\left[[\bm{\mathcal{C}}]_{(1)}\big(\overline{\bm{G}}\otimes\bm{\Psi}\otimes\bm{X}\big)^{\text{T}}\right]^{\dagger}$ (50)
    $\hat{\bm{X}}=[\bm{\mathcal{Q}}]_{(2)}\left[[\bm{\mathcal{C}}]_{(2)}\big(\overline{\bm{G}}\otimes\bm{\Psi}\otimes\bm{HS}\big)^{\text{T}}\right]^{\dagger}$ (51)
    $\hat{\overline{\bm{G}}}=[\bm{\mathcal{Q}}]_{(4)}\left[[\bm{\mathcal{C}}]_{(4)}\big(\bm{\Psi}\otimes\bm{X}\otimes\bm{HS}\big)^{\text{T}}\right]^{\dagger}.$ (52)
@@
    input : Received signal $\bm{\mathcal{Q}}$
    output : $\hat{\bm{H}}$, $\hat{\overline{\bm{G}}}$ and $\hat{\bm{X}}$
    1exto $i=0$ initialize $\hat{\overline{\bm{G}}}_{(i=0)}$ and $\hat{\bm{X}}_{(i=0)}$ begin
@@
      Find a least squares estimate of $\hat{\bm{HS}}_{i}$:
    • $\hat{(\bm{HS})}_{i}=[\bm{\mathcal{Q}}]_{(1)}\big[\underbrace{[\bm{\mathcal{C}}]_{(1)}\big(\overline{\bm{G}}_{(i-1)}\otimes\bm{\Psi}\otimes\bm{X}_{(i-1)}\big)^{\text{T}}}_{N\times IKT}\big]^{\dagger}$
@@
    Find a least squares estimate of $\hat{\bm{X}}_{i}$:
    • $\hat{\bm{X}}_{i}=[\bm{\mathcal{Q}}]_{(2)}\big[\underbrace{[\bm{\mathcal{C}}]_{(2)}\big(\overline{\bm{G}}_{(i-1)}\otimes\bm{\Psi}\otimes(\bm{HS})_{i}\big)^{\text{T}}}_{M_{T}\times IKM_{R}}\big]^{\dagger}$
@@
    Find a least squares estimate of $\hat{\overline{\bm{G}}}_{i}$:
    • $\hat{\overline{\bm{G}}}_{i}=[\bm{\mathcal{Q}}]_{(4)}\big[\underbrace{[\bm{\mathcal{C}}]_{(4)}\big(\bm{\Psi}\otimes\bm{X}_{i}\otimes(\bm{HS})_{i}\big)^{\text{T}}}_{M_{T}N\times KTM_{R}}\big]^{\dagger}$
@@
    Repeat steps $1$ to $3$ until convergence. end
@@
    From (45), an alternative (closed-form) approach can be used to jointly estimate the involved channels $\bm{H}$, $\bm{G}$ as well as the data symbol matrix $\bm{X}$. This approach left-filters the $3$-mode unfolding of the received signal using $(\bm{\Psi}[\bm{\mathcal{C}}]_{(1)})^{\dagger}$ followed by solving a “triple-Kronecker” factorization problem o the filtered signal. The algorithm is similar to the one proposed in [SOKAL_Elsevier]. Despite being a less complex estimation method, this approach imposes a more restrictive restriction on the data block length, requiring $K\geq(M_{T}N)^{2}$.